OpenStudy (johnw):

lim x->0 x⋅e^sin(x)/x Calculate the limit.

OpenStudy (eliesaab):

Is it \[ e^{\frac{\sin(x)}x} \]

OpenStudy (johnw):

Yup! :D

OpenStudy (eliesaab):

What is the limit of \[ \lim_{x\to0}\frac{\sin(x)}x \]

OpenStudy (johnw):

Are you asking me a different question or asking if this is my question?

OpenStudy (eliesaab):

That is part of your problem, so it is a different question that will help you decide about your limit

OpenStudy (eliesaab):

If the limit I asked you about is \(c\), then your limit will be \(e^c\)

OpenStudy (johnw):

well, sin(x)/x is 1 right?

OpenStudy (johnw):

Yes, an elementary limit was the word I was looking for :)

OpenStudy (eliesaab):

So your original limit is \[ e^1=e\]

OpenStudy (johnw):

I´m sorry, but I´ll need more step by step help.... :P

OpenStudy (eliesaab):

\[ \lim_{x\to 0} \, e^{\frac{\sin (x)}{x}}=e^{\lim_{x\to 0} \, \frac{\sin (x)}{x}}=e^1=e \]

OpenStudy (johnw):

But what happened to x in X⋅e^sin(x)/x?

OpenStudy (mathmate):

`is it ` \(e^{\frac{\sin(x)}x}\) `?` `yup! :D`.

OpenStudy (mathmate):

With the factor x, the limit becomes 0\(\times\)e=?

OpenStudy (johnw):

so in the end we´ll have 0 x e^1? Is the limit 0?

OpenStudy (mathmate):

Yes, with the multiplier x.

OpenStudy (johnw):

with the multiplier x? Doesn´t that make it automatically 0?

OpenStudy (mathmate):

Not all the time. Try \(f(x)=x\times \frac{1}{(1-x)^2-1}\) as x-> 0